OpenStudy (anonymous):

The height of a cylinder varies inversely with its radius. If the height is doubled, what is the effect on the volume of the cylinder?

OpenStudy (anonymous):

The height is the inverse of the radius: \[r = \frac{1}{h}\] \[V = \pi r^{2}h = \pi \left(\frac{1}{h} \right)^{2}h = \pi \frac{1}{h}\] So if you double the height:\[V = \pi \frac{1}{2h}\] The volume is halved. I'm pretty sure there's a more elegant way of doing this with calculus but this works too.

OpenStudy (anonymous):

r=k/h volume of cylinder= pi r^2 h pi (k/2h)^2 (2h) = pi (k/4h ) 2h = 1/4 *2 pi(k/h)^2 h = 1/2 r^2 h

OpenStudy (anonymous):

I dont understand. Why are you guys saying r = k/h, but its h = k/r, "height varies inversely with radius". Does it matter which you put it as?